# OODTE: A Differential Testing Engine for the ONNX Optimizer

### v3

@@ 10. Conclusion
We propose OODTE, a tool for the differential testing of the ONNX Optimizer. Using OODTE, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, OODTE identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.

### v5

@@ 12. Conclusion
We propose DiTOX, a tool for the differential testing of the ONNX Optimizer. Using DiTOX, we evaluated $130$ models across classification, object detection, and text generation tasks. We detected 9.29.2% of model instances that crashed when using the primary optimizer strategies, while 3030% of classification models and 16.616.6% of object detection models showed differences between the source and optimized versions, while text generation models remained robust. Additionally, DiTOX identified 15 issues, 14 of them new, related to crashes and accuracy degradation across 9 optimizer passes, which we reported to the ONNX Optimizer developers.